# Quadratic Modelings of Syndrome Decoding

### v1

@@ 6.3 The Dimension of the Variety Associated with Modeling 5
The parity-check matrix $\mathbf{H}$ has $i$ linearly dependent columns indexed by the set $S$ iff the corresponding code $\mathcal{C}$ has a codeword of weight $\leq i$ with support contained in $S$. Hence any codeword of positive weight $i\leq t$ with support $S^{\prime}$ identifies a set of $\binom{n-i}{t-i}$ supersets $S\supseteq S^{\prime}$. For each of such $S$, the inequality $\mathsf{rk}(H_{S})<i+(t-i)=t$ holds. If the syndrome $\mathbf{s}\in\mathbb{F}_{q}^{n-k}$ is sampled at random, then it belongs to the column space of $S^{\prime}$ with probability $\frac{1}{q^{n-k-i+1}}$. Otherwise, for any of the $\binom{n-i}{t-i}$ possible choices of $S$, the probability that $\mathbf{s}$ lies in the column space of $\mathbf{H}_{S}$ but not in that of $\mathbf{H}_{S^{\prime}}$ is $\frac{1}{q^{n-k-t+1}}-\frac{1}{q^{n-k-i+1}}$. Let $W_{i}(\mathcal{C})$ be the number of codewords in $\mathcal{C}$ of weight exactly $i$. By summing the probabilities of the two mutually disjoint events and from the proof of Proposition 5, we can thus upper bound the probability that the ideal is positive-dimensional for a random syndrome as
$\sum_{i=1}^{t}W_{i}(\mathcal{C})\left(\frac{1}{q^{n-k-i+1}}+\binom{n-i}{t-i}\left(\frac{1}{q^{n-k-t+1}}-\frac{1}{q^{n-k-i+1}}\right)\right).$
In the case of the codeword finding problem, i.e. if the syndrome is the zero vector, then the upper bound on the sought probability simplifies to
$\sum_{i=1}^{t}W_{i}(\mathcal{C})\binom{n-i}{t-i},$
since the zero vector syndrome belongs to any column space. ∎
@@
In Appendix 0.A, Tables 6 and 7, we provide examples of such bounds for concrete parameters. In the case of syndrome decoding, these probabilities are very small even at Gilbert-Varshamov distance and the issue of having ideals of positive dimension is thus absolutely negligible for the purpose of cryptanalysis. On the opposite, the bound on the probability of the same event for the codeword finding problem becomes completely useless when approaching the Gilbert-Varshamov distance. Indeed, the trivial upper bound “${\mathbb{P}}\leq 1$” entries from the two tables mean that the bound given by Proposition 6 gives a number larger than 1. A possible workaround for the described issue with the codeword finding version is to make use of hybrid methods. Indeed, it is enough to guess a number of nonzero positions equal or greater than the ideal dimension to decrease the latter to 0 with high probability. Recalling that the solution space is projective and one the value of one nonzero entry can be chosen arbitrarily, specializing $l$ coordinates has a success probability of $\frac{\binom{n-l}{t-l}}{\binom{n}{t}(q-1)^{l-1}}$. In cryptanalysis, however, it is usually assumed to know the minimal weight of a (nonzero) solution. This is because, if there exists a solution of weight smaller than the target, then the challenge is actually easier. The next corollary thus implies a strategy to obtain a zero-dimensional ideal in this setting. If we suppose to know a lower bound $d^{\prime}$ the minimum distance $d(C)$ of the code, then it means that any $d^{\prime}-1$ columns of the parity-check matrix $\mathbf{H}$ are linearly independent. Therefore, Equation (19) implies that the dimension of the ideal for the bounded weight modeling with target weight $d^{\prime}$ is exactly 1 and thus it is enough to specialize one variable $x_{i}$ to any element in $\mathbb{F}_{q}$ to obtain a zero-dimensional ideal.
@@
$n$ $k$ $t$Lin $r_{1}$ $r_{2}$Quad #Vars SD $r_{1}$ $r_{2}$Quad #Vars SD $r_{1}$ $r_{2}$Quad #Vars SD
@@
Table 5: This table gives information from experiments using random $\mathbb{F}_{q}$-linear codes using Modeling 5. The values in the SD column represent the highest step degree achieved when directly computing the Gröbner basis of the system in MAGMA. This is typically regarded as a proxy for the solving degree $d_{\mathrm{sol}}$. The column “#lin” denotes the number of linear equations, i.e. of parity-check equations, which is independent from the field size. The columns “#quad” and “#vars” stand for the number of quadratic equations and the number of variables, which depend on the value $r_{2}$ instead. The integer $r_{1}$ is the extension field degree over which the equations are defined. We recall that the value $m$ leads to different possible choices of $(r_{1},r_{2})$ and we give all minima with respect to the standard partial order on pairs.
@@ Acknowledgments.
$t=\lfloor(d_{GV}-1)/2\rfloor$ $t=\lfloor(n-k)/2\rfloor$ $t=d_{GV}$ $t=d_{GV}+1$
@@
$[1000,500]_{2}$ $t=55,$ ${\mathbb{P}}\leq 1.05\cdot 10^{-177}$ $t=250,$ ${\mathbb{P}}\leq 1$ $t=112,$ ${\mathbb{P}}\leq 9.96\cdot 10^{-84}$ $t=113,$ ${\mathbb{P}}\leq 3.13\cdot 10^{-82}$
$[1000,500]_{7}$ $t=119,$ ${\mathbb{P}}\leq 1.85\cdot 10^{-488}$ $t=250,$ ${\mathbb{P}}\leq 1.95\cdot 10^{-181}$ $t=239,$ ${\mathbb{P}}\leq 2.06\cdot 10^{-205}$ $t=240,$ ${\mathbb{P}}\leq 3.20\cdot 10^{-203}$
$[1000,500]_{127}$ $t=182,$ ${\mathbb{P}}\leq 2.66\cdot 10^{-1136}$ $t=250,$ ${\mathbb{P}}\leq 4.75\cdot 10^{-812}$ $t=366,$ ${\mathbb{P}}\leq 5.44\cdot 10^{-283}$ $t=367,$ ${\mathbb{P}}\leq 1.52\cdot 10^{-278}$
$[1000,800]_{2}$ $t=16,$ ${\mathbb{P}}\leq 1.41\cdot 10^{-79}$ $t=100,$ ${\mathbb{P}}\leq 1$ $t=32,$ ${\mathbb{P}}\leq 8.22\cdot 10^{-42}$ $t=33,$ ${\mathbb{P}}\leq 9.65\cdot 10^{-40}$
$[1000,800]_{7}$ $t=36,$ ${\mathbb{P}}\leq 1.30\cdot 10^{-212}$ $t=100,$ ${\mathbb{P}}\leq 8.72\cdot 10^{-31}$ $t=74,$ ${\mathbb{P}}\leq 2.94\cdot 10^{-101}$ $t=75,$ ${\mathbb{P}}\leq 1.78\cdot 10^{-98}$
$[1000,800]_{127}$ $t=61,$ ${\mathbb{P}}\leq 3.32\cdot 10^{-489}$ $t=100,$ ${\mathbb{P}}\leq 8.70\cdot 10^{-284}$ $t=124,$ ${\mathbb{P}}\leq 3.03\cdot 10^{-161}$ $t=125,$ ${\mathbb{P}}\leq 3.43\cdot 10^{-156}$
@@
$t=\lfloor(d_{GV}-1)/2\rfloor$ $t=\lfloor(n-k)/2\rfloor$ $t=d_{GV}-2$ $t=d_{GV}-1$ $t=d_{GV}$
@@
$[1000,500]_{2}$ $t=55,$ ${\mathbb{P}}\leq 1.91\cdot 10^{-43}$ $t=250,$ ${\mathbb{P}}\leq 1$ $t=110,$ ${\mathbb{P}}\leq 1$ $t=111,$ ${\mathbb{P}}\leq 1$ $t=112,$ ${\mathbb{P}}\leq 1$
$[1000,500]_{7}$ $t=120,$ ${\mathbb{P}}\leq 1.24\cdot 10^{-165}$ $t=250,$ ${\mathbb{P}}\leq 1$ $t=237,$ ${\mathbb{P}}\leq 1$ $t=238,$ ${\mathbb{P}}\leq 1$ $t=239,$ ${\mathbb{P}}\leq 1$
$[1000,500]_{127}$ $t=182,$ ${\mathbb{P}}\leq 3.46\cdot 10^{-463}$ $t=250,$ ${\mathbb{P}}\leq 5.40\cdot 10^{-284}$ $t=364,$ ${\mathbb{P}}\leq 0.00115$ $t=365,$ ${\mathbb{P}}\leq 0.255$ $t=366,$ ${\mathbb{P}}\leq 1$
$[1000,800]_{2}$ $t=15,$ ${\mathbb{P}}\leq 1.40\cdot 10^{-23}$ $t=100,$ ${\mathbb{P}}\leq 1$ $t=30,$ ${\mathbb{P}}\leq 1$ $t=31,$ ${\mathbb{P}}\leq 1$ $t=32,$ ${\mathbb{P}}\leq 1$
$[1000,800]_{7}$ $t=36,$ ${\mathbb{P}}\leq 3.60\cdot 10^{-73}$ $t=100,$ ${\mathbb{P}}\leq 1$ $t=72,$ ${\mathbb{P}}\leq 1$ $t=73,$ ${\mathbb{P}}\leq 1$ $t=74,$ ${\mathbb{P}}\leq 1$
$[1000,800]_{127}$ $t=61,$ ${\mathbb{P}}\leq 1.13\cdot 10^{-194}$ $t=100,$ ${\mathbb{P}}\leq 2.66\cdot 10^{-71}$ $t=122,$ ${\mathbb{P}}\leq 3.67\cdot 10^{-5}$ $t=123,$ ${\mathbb{P}}\leq 0.0333$ $t=124,$ ${\mathbb{P}}\leq 1$

### v2

@@ 6.3 The Dimension of the Variety Associated with Modeling 5
It follows from Corollary 3 that the variety has positive dimension if and only if there exists a set $S\subseteq[n]$, $|S|=t$, such that $\mathbf{H}_{S}$ is not full-rank and the syndrome $\mathbf{s}$ belongs to the column space of $\mathbf{H}_{S}$.
The parity-check matrix $\mathbf{H}$ has $i$ linearly dependent columns indexed by the set $S$ if and only if the corresponding code $\mathcal{C}$ has a codeword of weight $\leq i$ with support contained in $S$. Hence any codeword of positive weight $i\leq t$ with support $S^{\prime}$ identifies a set of $\binom{n-i}{t-i}$ supersets $S\supseteq S^{\prime}$. On the other hand, each set $S$ of $t$ dependent columns is associated with at least one codeword of weight $\leq t$, hence iterating over such codewords is enough to guarantee an upper bound.
Let $W_{i}(\mathcal{C})$ be the number of codewords in $\mathcal{C}$ of weight exactly $i$. By splitting the event $\{\mathbf{s}\in\mathsf{ColSpace}(\mathbf{H}_{S})\}$ into the union of the two disjoint events $\{\mathbf{s}\in\mathsf{ColSpace}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})\}$ and $\{\mathbf{s}\in\mathsf{ColSpace}(\mathbf{H}_{S})\setminus\mathsf{ColSpace}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})\}$ and using that $\mathsf{rk}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})\leq i-1$ and $\mathsf{rk}(\mathbf{H}_{S})\leq\mathsf{rk}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})+(t-i)$, we thus obtain an upper bound on the sought probability:
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}S\subseteq[n]\\ |S|=t\land\mathsf{rk}(\mathbf{H}_{S})<t\end{subarray}}\{\mathbf{s}\in\mathsf{ColSpace}(\mathbf{H}_{S})\}\right)$
$\leq$ $\sum_{\begin{subarray}{c}S\subseteq[n]\\ |S|=t\land\mathsf{rk}(\mathbf{H}_{S})<t\end{subarray}}\mathbb{P}(\{\mathbf{s}\in\mathsf{ColSpace}(\mathbf{H}_{S})\})$
$\leq$ $\sum_{\begin{subarray}{c}\mathbf{c}\in\mathcal{C}\\ \mathsf{wt}(\mathbf{c})\leq t\end{subarray}}\left(\mathbb{P}(\{\mathbf{s}\in\mathsf{ColSpace}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})\})+\sum_{\begin{subarray}{c}\mathsf{supp}(\mathbf{c})\subseteq S\subseteq[n]\\ |S|=t\end{subarray}}\mathbb{P}(\{\mathbf{s}\in\mathsf{ColSpace}(\mathbf{H}_{S})\setminus\mathsf{ColSpace}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})\})\right)$
$=$ $\sum_{i=1}^{t}W_{i}(\mathcal{C})\left(\frac{q^{\dim_{\mathbb{F}_{q}}\mathsf{ColSpace}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})}}{q^{n-k}}+\binom{n-i}{t-i}\frac{q^{\dim_{\mathbb{F}_{q}}\mathsf{ColSpace}(\mathbf{H}_{S})}-q^{\dim_{\mathbb{F}_{q}}\mathsf{ColSpace}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})}}{q^{n-k}}\right)$
$=$ $\sum_{i=1}^{t}W_{i}(\mathcal{C})\left(\frac{q^{\mathsf{rk}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})}}{q^{n-k}}+\binom{n-i}{t-i}\frac{q^{\mathsf{rk}(\mathbf{H}_{S})}-q^{\mathsf{rk}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})}}{q^{n-k}}\right)$
$\leq$ $\sum_{i=1}^{t}W_{i}(\mathcal{C})\left(\frac{q^{\mathsf{rk}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})}}{q^{n-k}}+\binom{n-i}{t-i}\frac{q^{\mathsf{rk}(\mathbf{H}_{\mathsf{supp}(\mathbf{c})})}}{q^{n-k}}(q^{t-i}-1)\right)$
$\leq$ $\sum_{i=1}^{t}W_{i}(\mathcal{C})\left(\frac{1}{q^{n-k-i+1}}+\binom{n-i}{t-i}\left(\frac{1}{q^{n-k-t+1}}-\frac{1}{q^{n-k-i+1}}\right)\right).$
Finally, in the case of the codeword finding problem, i.e. if the syndrome is the zero vector, the condition on the positive dimension of the variety boils down to the existence of the set $S$, $|S|=t$, such that the rank $\mathbf{H}_{S}$ is defective, as the zero vector belongs to any linear subspace. Therefore, in this case, the calculations are simplified into:
$\mathbb{P}\left(\bigcup_{\begin{subarray}{c}S\subseteq[n]\\ |S|=t\land\mathsf{rk}(\mathbf{H}_{S})<t\end{subarray}}\{0\in\mathsf{ColSpace}(\mathbf{H}_{S})\}\right)\leq\sum_{\begin{subarray}{c}S\subseteq[n]\\ |S|=t\land\mathsf{rk}(\mathbf{H}_{S})<t\end{subarray}}1\leq\sum_{i=1}^{t}W_{i}(\mathcal{C})\binom{n-i}{t-i}.$
@@
In Appendix 0.A, Tables 6 and 7, we provide examples of such bounds for concrete parameters. In the case of syndrome decoding, these probabilities are very small even at Gilbert-Varshamov distance and the issue of having ideals of positive dimension is thus absolutely negligible for the purpose of cryptanalysis. On the opposite, the bound on the probability of the same event for the codeword finding problem becomes completely useless when approaching the Gilbert-Varshamov distance. Indeed, the trivial upper bound “${\mathbb{P}}\leq 1$” entries from the two tables mean that the bound given by Proposition 6 gives a number larger than 1. A possible workaround for the described issue with the codeword finding version is to make use of hybrid methods. Indeed, it is enough to guess a number of nonzero positions equal or greater than the ideal dimension to decrease the latter to 0 with high probability. Recalling that the solution space is projective and one the value of one nonzero entry can be chosen arbitrarily, specializing $l$ coordinates has a success probability of $\frac{\binom{n-l}{t-l}}{\binom{n}{t}(q-1)^{l-1}}$. In cryptanalysis, however, it is usually assumed to know the minimal weight of a (nonzero) solution. This is because, if there exists a solution of weight smaller than the target, then the challenge is actually easier.
A simple strategy to obtain a zero-dimensional ideal in this setting is thus the following. If we suppose to know a lower bound $d^{\prime}$ the minimum distance $d(C)$ of the code, then it means that any $d^{\prime}-1$ columns of the parity-check matrix $\mathbf{H}$ are linearly independent. Therefore, Equation (19) implies that the dimension of the ideal for the bounded weight modeling with target weight $d^{\prime}$ is exactly 1 and thus it is enough to specialize one variable $x_{i}$ to any element in $\mathbb{F}_{q}$ to obtain a zero-dimensional ideal.
@@
$n$ $k$ $t$lin $r_{1}$ $r_{2}$quad #vars $d_{\mathrm{M}}$ $r_{1}$ $r_{2}$quad #vars $d_{\mathrm{M}}$ $r_{1}$ $r_{2}$quad #vars $d_{\mathrm{M}}$
@@
Table 5: This table gives information from experiments using random $\mathbb{F}_{q}$-linear codes using Modeling 5. The values in the $d_{\mathrm{M}}$ column represent the highest step degree achieved when directly computing the Gröbner basis of the system in MAGMA. The column “#lin” denotes the number of linear equations, i.e. of parity-check equations, which is independent from the field size. The columns “#quad” and “#vars” stand for the number of quadratic equations and the number of variables, which depend on the value $r_{2}$ instead. The integer $r_{1}$ is the extension field degree over which the equations are defined. We recall that the value $m$ leads to different possible choices of $(r_{1},r_{2})$ and we give all minima with respect to the standard partial order on pairs.
@@ Acknowledgments.
We would like to thank the reviewers for their detailed and valuable feedback. Additionally, special thanks to Magali Bardet, Tanja Lange and Alberto Ravagnani for the fruitful discussions and insights.
@@
$[n,k]_{q}$ $t=\lfloor(d_{GV}-1)/2\rfloor$ $t=\lfloor(n-k)/2\rfloor$ $t=d_{GV}$ $t=d_{GV}+1$
@@
$[n,k]_{q}$ $t=\lfloor(d_{GV}-1)/2\rfloor$ $t=\lfloor(n-k)/2\rfloor$ $t=d_{GV}-2$ $t=d_{GV}-1$ $t=d_{GV}$